# Denoising Diffusion Probabilistic Models for Coastal Inundation Forecasting

### v1

@@ 5 Diff-Flood Method
As described in Section B, within the denoising diffusion model, we need to train a neural network $\hat{\mathbf{x}}_{\theta_{3}}({}_{k}\mathbf{x}^{n}_{t},n,{}_{k}\mathbf{h}_{t_{0}:t-1})$ that predicts ${}_{k}\mathbf{x}^{0}_{t}$ from noisy vector ${}_{k}\mathbf{x}^{n}_{t}$, noise step $n$, and context embedding ${}_{k}\mathbf{h}_{t_{0}:t-1}$. We use a conditional UNet for this purpose. Within the UNet architecture, we employ a cross-attention conditioning mechanism (Rombach et al., 2022) to map the context embedding to the intermediate layers of the UNet. It allows us to predict ${}_{k}\mathbf{x}^{0}_{t}$ while paying attention to different parts of the context embedding.
5.1 Training
Input: Data: ${}_{k}\mathbf{x}^{0}_{t}\sim q({}_{k}\mathbf{x}^{0}_{t})$, Context: ${}_{k}{\mathbf{x}^{0}_{t_{0}:t-1}}$, ${}_{k}\mathbf{z}_{t_{0}:t-1}$, ${}_{k}\mathbf{s}$.
@@ 5.1 Training
       // Compute context embedding
4       ${}_{k}\mathbf{f}^{0}_{t_{0}:t-1}\leftarrow\textsc{conv}{}_{\theta_{1}}({}_{k}\mathbf{x}^{0}_{t_{0}:t-1}\oplus{}_{k}\mathbf{s})$
5       ${}_{k}\mathbf{g}^{0}_{t_{0}:t-1}\leftarrow\textsc{sinusoid}{}({}_{k}\mathbf{z}_{t_{0}:t-1})$ ${}_{k}\mathbf{h}_{t_{0}:t-1}\leftarrow\textsc{linear}{}_{\theta_{2}}({}_{k}\mathbf{f}_{t_{0}:t-1}\oplus{}_{k}\mathbf{g}_{t_{0}:t-1})$
@@
7       Take gradient step on $\nabla_{\theta}||{}_{k}\mathbf{x}^{0}_{t}-{}_{k}\hat{\mathbf{x}}^{0}_{t}||^{2}_{2}$
Algorithm 1 Training for patch $P_{k}$, prediction time step $t$
To train Diff-Flood, we sample context windows (${}_{k}\mathbf{x}^{0}_{t_{0}:t-1}$) and the immediate next timestep (${}_{k}\mathbf{x}^{0}_{t}$) from the training time series data of each patch. We then optimize the parameters $\theta$ by minimizing the negative log-likelihood of the data ($-\log p_{\theta}({}_{k}\mathbf{x}^{0}_{t}|{}_{k}\mathbf{h}_{t_{0}:t-1}$). Through a similar derivation as shown in the Section B, the objective to minimize becomes:
$\mathbb{E}_{{}_{k}\mathbf{x}^{0}_{t},n}||{}_{k}\mathbf{x}^{0}_{t}-\hat{\mathbf{x}}_{\theta}({}_{k}\mathbf{x}_{t}^{n},n,{}_{k}\mathbf{h}_{t_{0}:t-1}||^{2}_{2}$ (15)
Here, the neural network $\hat{x}_{\theta}$ is now also conditioned on the context embedding ${}_{k}\mathbf{h}_{t_{0}:t-1}$.
Input: Noise: ${}_{k}\mathbf{x}^{N}_{t}\sim\mathcal{N}({\mathbf{0}},\mathbf{I})$, Context: ${}_{k}{\mathbf{x}^{0}_{t_{0}:t-1}}$, ${}_{k}\mathbf{z}_{t_{0}:t-1}$, ${}_{k}\mathbf{s}$.
@@
3             $\mathbf{v}\sim\mathcal{N}({\mathbf{0}},\mathbf{I})$
4      else
5             $\mathbf{v}\leftarrow{\mathbf{0}}$
@@
9       $\hskip 160.00024pt\ignorespaces+\sqrt{\Sigma_{\theta}}\mathbf{v}$
@@
5.2 Inference
@@
During inference, given a context ${}_{k}\mathbf{x}^{0}_{t_{0}:t-1},{}_{k}\mathbf{z}_{t_{0}:t-1},{}_{k}\mathbf{s}$, we first compute the context embedding ${}_{k}\mathbf{h}_{t_{0}:t-1}$. We then apply Algorithm 2 to obtain a sample ${}_{k}\mathbf{x}^{0}_{t}$ of the next time step. Based on this sample, we compute the new context embedding ${}_{k}\mathbf{h}_{t_{0}:t}$ and repeat until the desired number of prediction time steps have been reached. We repeat the process multiple times to get multiple predictions.
6 Experiment
@@
We conduct our experiments in two parts. First, we evaluate Diff-Flood model on the TideWatch dataset (see Section 6.1) against five competitive baseline models, taking forecasting inundation across the Virginian Eastern Shore region as our case study. Second, we conduct an ablation study of our model, underscoring the importance of the components of our proposed architecture.
6.1 Dataset
@@
To train Diff-Flood, we downloaded inundation data of the Eastern Shore, Virginia from Tidewatch Maps (Loftis et al., 2019; Loftis, 2022); a street-level inundation mapping tool developed by the Virginia Institute of Marine Science (VIMS). It utilizes the open-source hydro-dynamic model, SCHISM (Zhang et al., 2016) as the engine, and provides 36 hour inundation forecast maps with a 12 hour update frequency. We used this dataset due to its high spatial resolution, coverage of areas affected by coastal inundation, and lack of other publicly available inundation datasets.
@@
We downloaded 89 days (March 28, 2024 to June 24, 2024) of hourly inundation data of the area. The data comes in a raster geo-database format with a pixel resolution of $10m\times 10m$ and provides inundation level at each pixel in feet. We resampled the data to a pixel resolution of $30m\times 30m$. We then consider different square-shaped patches of size: $16\times 16$, $64\times 64$, $80\times 80$, and $96\times 96$. Each patch $P_{k}$ has a time series of inundation levels denoted by ${}_{k}\mathbf{x}^{0}_{t}\in\mathbb{R}^{D\times D}$, where $D$ is the length of one side of the patch. We aim to train Diff-Flood to forecast inundation levels on multiple patches.
We have used 81 days (1944 timesteps, each of length 1-hour) of data for training, 1 day (24 timesteps) of data for validation, and 7 days (168 timesteps) of data for testing. Note that, when we utilize $K$ number of patches, the total number of both training and testing timesteps are multiplied by $K$. In our forecasting task, we use context window of $12$ hours, and then predict the inundation levels in the subsequent $12$ hours (prediction window). Each sample datapoint therefore contains 24 consecutive timesteps, first $12$ hours for context and the last $12$ hours for prediction. We normalize each datapoint separately as follows: first we calculate the mean and standard deviation of the inundation values in the context window, then we standardize the inundation values in both the context and prediction window using the calculated mean and standard deviation. In addition to the inundation data, we used digital elevation data (OpenTopography, 2021) for the same area. We resampled the elevation data to the same resolution of $30m\times 30m$ and then aligned it with our inundation data.
6.2 Evaluation Metric
We use two metrics to evaluate forecast quality with respect to the observations. The first metric is Normalized Root Mean Squared Error (NRMSE). Let, $\mathbf{x}$ be the set of all observations over all cells for all timesteps. $x_{ij,t}\in\mathbf{x}$ be the observed value of cell $(i,j)$ at time $t$ and let, $\tilde{x}_{ij,t}$ denote the corresponding average forecast value. Then, NRMSE is defined as follows:
In addition to NRMSE, we use an extended version of the Continuous Ranked Probability Score (CRPS) Winkler et al. (1996) to quantify forecast uncertainty. The CRPS metric is calculated for each cell and each timestep individually. Let’s assume that we have $M$ forecasts ${\hat{x}_{ij,t}^{1},\hat{x}_{ij,t}^{2},...,\hat{x}_{ij,t}^{M}}$ for observation $x_{ij,t}$. First, the empirical cumulative distribution function (CDF) at a point $z$, $\hat{F}_{ij,t}^{M}(z)$ from these forecast values is defined as
where $\mathbb{I}(y)$ is the indicator function that yields $1$ if condition $y$ holds, 00 otherwise. From this, CRPS between the empirical CDF $\hat{F}_{ij,t}^{M}$ and the observation $x_{ij,t}$ is calculated as
@@ 6.2 Evaluation Metric
6.3 Baseline Comparison
Patch
$D\times D\times K$
Metric DiffSTG DCRNN DeepVAR LSTNet TimeGrad Diff-Flood
$16\times 16\times 2$ NACRPS 1.2523$\pm$0.0445 N/A 1.4782$\pm$0.0332 N/A 1.2319$\pm$0.0672 0.9603$\pm$0.2119 ($\uparrow$22.05%)
NRMSE 0.1278$\pm$0.0022 0.1302$\pm$0.0125 0.1742$\pm$0.0034 0.1439$\pm$0.0003 0.1601$\pm$0.0071 0.1162$\pm$0.0189 ($\uparrow$9.02%)
$16\times 16\times 5$ NACRPS 1.3166$\pm$0.0164 N/A 1.4151$\pm$0.0293 N/A 1.2558$\pm$0.0707 0.834$\pm$0.1781 ($\uparrow$33.59%)
NRMSE 0.1373$\pm$0.0015 0.1166$\pm$0.0096 0.1734$\pm$0.0023 0.1501$\pm$0.0001 0.17$\pm$0.0075 0.1095$\pm$0.0165 ($\uparrow$6.13%)
$64\times 64\times 10$ NACRPS Failed N/A 1.0228$\pm$0.0895 N/A 0.5522$\pm$0.0912 0.2007$\pm$0.0201 ($\uparrow$63.66%)
NRMSE 0.0801$\pm$0.0065 0.3437$\pm$0.035 0.1765$\pm$0.0002 0.1818$\pm$0.0182 0.0639$\pm$0.0079 ($\uparrow$20.21%)
$64\times 64\times 20$ NACRPS Failed Failed 1.1451$\pm$0.0792 N/A 0.545$\pm$0.0181 0.227$\pm$0.0249 ($\uparrow$58.34%)
NRMSE 0.3967$\pm$0.0419 0.1941$\pm$0.0006 0.1908$\pm$0.0029 0.0794$\pm$0.008 ($\uparrow$58.39%)
$80\times 80\times 20$ NACRPS Failed Failed 1.1494$\pm$0.1409 N/A 0.5251$\pm$0.0105 0.2091$\pm$0.0216 ($\uparrow$60.18%)
NRMSE 0.2761$\pm$0.0399 0.131$\pm$0.0003 0.1251$\pm$0.0021 0.0508$\pm$0.0034 ($\uparrow$59.39%)
$96\times 96\times 20$ NACRPS Failed Failed 1.3578$\pm$0.0372 N/A 0.5788$\pm$0.016 0.2519$\pm$0.0241 ($\uparrow$56.48%)
NRMSE 0.1661$\pm$0.0057 0.0748$\pm$0.0001 0.0718$\pm$0.001 0.0334$\pm$0.0044 ($\uparrow$53.44%)
Table 1: Performance Comparison of Diff-Flood with baseline models. All metrics are reported up to one standard deviation after conducting five experiments with different random seeds. For each configuration, the best performance is bold-highlighted and the second-best performance is underlined. For Diff-Flood, percentage improvement in performance with its closest baseline competitor is shown. State-of-the-art spatiotemporal forecasting methods (e.g., DiffSTG, DCRNN) could not handle training for large patch configurations, these are marked as ‘Failed’. Since, DCRNN and LSTNet are point forecasting methods, their NACRPS metric are marked as ‘N/A’.
In this section, we describe the baseline models compared with our model. All experiments were conducted on an HPC cluster. To achieve scalability, experiments with different configurations were conducted with the SLURM scheduler. Each job used up to $256$GB CPU RAM, $8$ CPU cores and could request for one GPU with maximum $40$ GB GPU RAM.
@@ 6.3 Baseline Comparison
    DiffSTG (Wen et al., 2023): The State-of-the-art (SOTA) diffusion-based model for spatiotemporal forecasting.
@@
    TimeGrad (Rasul et al., 2021) The SOTA diffusion-based model designed for Multivariate time-series forecasting.
@@
    DeepVAR (Salinas et al., 2019): Combines an RNN-based architecture with a Gaussian copula process output model with a low-rank covariance structure.
@@
    LSTNet (Lai et al., 2018): The Long-and Short-term Time-series network (LSTNet) captures local dependency pattern among variables and long-term temporal correlation among time-series trends using CNN and RNN.
@@
    DCRNN (Li et al., 2017): A spatiotemporal forecasting framework that combines Diffusion Convolution and Sequence-to-sequence architecture to capture temporal and spatial dependencies.
@@
Each model was trained on six different patch settings as shown in Table 1. All hyper-parameters of our model are provided in the supplementary material. It is to be noted that the number of variables increases quadratically with patch size and linearly with the number of patches. Since some models do not scale well [^1] It is worth mentioning we also attempted to use $D^{3}$ (Li et al., 2022) as a baseline. However, we it scaled poorly; causing memory error even for our lowest patch configuration. with increments in the number of variables, we limit $D=96$ for this set of experiments. Across all experiments, we set the context and prediction length to 12. Eight scenarios were sampled to understand the quality of probabilistic forecasting. Table 1 summarizes the performance of our model with respect to the baseline models. Below, we outline some key observations from the results.
Figure 3: Scaling in Model Parameters (Log scale in Y axis) with increasing patch configuration. It has to be noted that model parameter is not affected by $K$. DiffSTG and DCRNN are not shown here since they fail to execute for large patch setting.
First, in terms of NACRPS, our model consistently outperforms the baselines, underscoring its superiority in probabilistic forecasting. For small patch size ($D=16$), Diff-Flood has a performance gain of between $22.05\%$ and $33.59\%$, compared to its closest baseline. For larger patch sizes ($D\geq 64$), the improvement in NACRPS is close to $60\%$, underscoring the performance scalability of our model. Second, Diff-Flood outperforms other baselines in terms of NRMSE too. Although TimeGrad was constantly the closest competitor baseline in terms of NACRPS, this is not the case here. For the smallest patch setting, DiffSTG was the best baseline in terms of NRMSE, which was outperformed by Diff-Flood by $9.02\%$. For the next two patch settings, DCRNN was the best baseline. In these cases, Diff-Flood outperformed DCRNN by $6.13\%$ and $20.21\%$, respectively. For the settings with $K=20$, DiffSTG and DCRNN both required extensive GPU memory due to poor scalability and they failed to train. In these settings, Diff-Flood outperformed TimeGrad with a margin of around $50-60\%$.

### v2

@@ Introduction
As described in the previous section, within the denoising diffusion model, we need to train a neural network $\hat{\mathbf{x}}_{\theta_{3}}({}_{k}\mathbf{x}^{n}_{t},n,{}_{k}\mathbf{h}_{t_{0}:t-1})$ that predicts $_{k}\mathbf{x}^{0}_{t}$ from noisy vector $_{k}\mathbf{x}^{n}_{t}$, noise step $n$, and context embedding $_{k}\mathbf{h}_{t_{0}:t-1}$. We use a conditional UNet for this purpose. Within the UNet architecture, we employ a cross-attention conditioning mechanism (Rombach_2022_CVPR) to map the context embedding to the intermediate layers of the UNet. It allows us to predict $_{k}\mathbf{x}^{0}_{t}$ while paying attention to the context embedding.
Training
Input: Data: $_{k}\mathbf{x}^{0}_{t}\sim q({}_{k}\mathbf{x}^{0}_{t})$, Context: $_{k}{\mathbf{x}^{0}_{t_{0}:t-1}}$, $_{k}\mathbf{z}_{t_{0}:t-1}$, $\mathbf{s}_{k}$.
@@
 // Masking
2   Randomly generate a binary sensor mask $\mathbf{M}_{k}$.
3 for $\tau\in[t_{0},t-1]$ do
4    $\mathbf{u}\sim\mathcal{N}({\mathbf{0}},\mathbf{I})$
5    $_{k}\mathbf{y}^{0}_{\tau}={}_{k}\mathbf{x}^{0}_{\tau}\odot\mathbf{M}_{k}+({\mathbf{1}}-\mathbf{M}_{k})\odot\mathbf{u}$
 // Compute context embedding
[…]
7 $_{k}\mathbf{g}_{t_{0}:t-1}\leftarrow\textsc{sinusoid}{}({}_{k}\mathbf{z}_{t_{0}:t-1})$
[…]
@@ 7 $_{k}\mathbf{g}_{t_{0}:t-1}\leftarrow\textsc{sinusoid}{}({}_{k}\mathbf{z}_{t_{0}:t-1})$
[…]
Algorithm 1 Training step on a data point of $P_{k}$.
Input: Noise: $_{k}\mathbf{x}^{N}_{t}\sim\mathcal{N}({\mathbf{0}},\mathbf{I})$, Context: $_{k}{\mathbf{y}^{0}_{t_{0}:t-1}}$, $_{k}\mathbf{z}_{t_{0}:t-1}$, $\mathbf{s}_{k}$, and $\mathbf{M}_{k}$.
@@ 2 $\mathbf{u}\leftarrow{\mathbf{0}}$
4    $\mathbf{u}\sim\mathcal{N}({\mathbf{0}},\mathbf{I})$
@@
8 $\hskip 160.00024pt\ignorespaces+\sqrt{\Sigma_{\theta}}\mathbf{u}$
@@
To train Diff-Sparse, we sample the following from the training time-series data of each patch: ($i$) inundation history in the context window ($_{k}\mathbf{x}^{0}_{t_{0}:t-1}$) and temporal co-variates ($_{k}\mathbf{z}_{t_{0},t-1}$), (iiii) inundation values in the immediate next timestep ($_{k}\mathbf{x}^{0}_{t}$). Diff-Sparse is trained to forecast the latter. We apply Algorithm 1 on each sample. First, we generate a random binary sensor mask $\mathbf{M}_{k}$. This mask is applied on the context inundation history ($_{k}\mathbf{x}^{0}_{t_{0}:t-1}$) to generate a sparse inundation history $_{k}\mathbf{y}^{0}_{t_{0}:t-1}$ (line 1-1, $\odot$ denotes element-wise multiplication). Here, we retain the inundation values in cells where sensors are placed (according to the sensor mask $\mathbf{M}_{k}$). For cells that do not have sensors, their inundation values are replaced with standard gaussian noise. This serves as a signal to the model to ignore these values. As we train Diff-Sparse with different random sensor masks, it is able to make forecasts based on different sensor placement configurations during inference. We use the sparse inundation history, temporal co-variates and the sensor mask to calculate the context embedding $_{k}\mathbf{h}_{t_{0}:t-1}$ (lines 1-1). We then optimize the model parameters $(\theta_{1},\theta_{2},\theta_{3})$ by minimizing the negative log-likelihood of the data conditioned on the context embedding ($\log p_{\theta}({}_{k}\mathbf{x}^{0}_{t}|{}_{k}\mathbf{h}_{t_{0}:t-1}$). Through a similar derivation as shown in the previous section, the objective to minimize becomes:
$\mathbb{E}_{{}_{k}\mathbf{x}^{0}_{t},n}||{}_{k}\mathbf{x}^{0}_{t}-\hat{\mathbf{x}}_{\theta_{3}}({}_{k}\mathbf{x}_{t}^{n},n,{}_{k}\mathbf{h}_{t_{0}:t-1}||^{2}_{2}$ (14)
Here, the neural network $\hat{x}_{\theta_{3}}$ is now also conditioned on the context embedding $_{k}\mathbf{h}_{t_{0}:t-1}$.
Inference
@@
During inference, given a context ($_{k}\mathbf{y}^{0}_{t_{0}:t-1}$, $_{k}\mathbf{z}_{t_{0}:t-1}$, $\mathbf{s}_{k}$, $\mathbf{M}_{k}$), we first compute the context embedding $_{k}\mathbf{h}_{t_{0}:t-1}$. We then apply Algorithm 2 to obtain a sample $_{k}\mathbf{x}^{0}_{t}$ of the next time step. Based on this sample, we compute the new context embedding $_{k}\mathbf{h}_{t_{0}:t}$ and repeat until the desired number of prediction time steps have been reached. We repeat the process multiple times to get multiple predictions.
Experiment
@@
In our experiments, we first evaluate Diff-Sparse on the TideWatch dataset against six competitive baseline models; taking forecasting inundation within the Virginia Eastern Shore region as our case study. Then, we conduct an ablation study to demonstrate the importance of the different context information utilized in Diff-Sparse.
Dataset
@@
To train Diff-Sparse, we downloaded inundation data of the Eastern Shore, Virginia from Tidewatch Maps (Loftis2019tidewatch; Loftis2022tidewatch); a street-level inundation mapping tool developed by the Virginia Institute of Marine Science (VIMS). It utilizes the open-source hydro-dynamic model, SCHISM (zhang2016seamless) as the engine, and provides 36 hour inundation forecast maps with a 12 hour update frequency. We used this dataset due to its high spatial resolution, coverage of areas affected by coastal inundation, and lack of other publicly available inundation datasets. Moreover, using it as training data allows Diff-Sparse to learn to make forecasts from a physics-based hydro-dynamic model. The dataset will be shared upon request.
@@
We downloaded 89 days (March 28, 2024 to June 24, 2024) of hourly inundation data of our study area. The data comes in a raster geo-database format with a pixel resolution of $10m\times 10m$ and provides inundation level at each pixel in feet. We resampled the data to a pixel resolution of $30m\times 30m$. We then consider different square-shaped patches. Each patch $P_{k}$ has a time series of inundation levels denoted by $_{k}\mathbf{x}^{0}_{t}\in\mathbb{R}^{D\times D}$, where $D\in\{16,64,80,96\}$ is the side length of the patch. We also used digital elevation data (DigitalElevation2021) for the same area. We resampled the elevation data to the same resolution and then aligned it with our inundation data. It serves as spatial context (i.e., elevation of patch $P_{k}$ denoted as $\mathbf{s}_{k}$) within Diff-Sparse.
We have used 81 days of data for training, 1 day of data for validation, and 7 days of data for testing. Training data points are sampled by following the process described in Section ‘Diff-Sparse Method (Training)’. We use a context window of $12$ hours. During testing, we predict the inundation levels in the subsequent $12$ hours. Inundation values are standardized based on the mean and standard deviation of the inundation values in the training data. Elevation values are standardized separately based on the mean and standard deviation of the elevation values in the study area.
Patch
$D^{2}\times K$
Metric DiffSTG DCRNN DeepVAR LSTNet TimeGrad BayesNF Diff-Sparse
@@
[…]
Table 1: Performance Comparison of Diff-Sparse with baseline models at 95% sparsity level (i.e., only $5\%$ of cells in a patch have sensors). All metrics are reported up to one standard deviation after conducting ten experiments with different random seeds. For each configuration, the best performance is bold-highlighted and the second-best performance is underlined. For Diff-Sparse, percentage improvement in performance with its closest baseline competitor is shown. State-of-the-art spatiotemporal forecasting methods (e.g., DiffSTG, DCRNN) could not handle training for large patch configurations; these are marked as ‘Failed’. Since, DCRNN and LSTNet are point forecasting methods, their NACRPS metrics are marked as ‘N/A’.
Evaluation Metric
We use two metrics to evaluate forecast quality. The first metric is Normalized Root Mean Squared Error (NRMSE). Let $\mathbf{x}$ be the set of all observations over all cells for all timesteps. Let $x_{ij,t}\in\mathbf{x}$ denote the observed inundation value of cell $(i,j)$ at time $t$ and let, $\tilde{x}_{ij,t}$ denote the corresponding average forecast value. Then:
$\text{NRMSE}=\tfrac{1}{\max_{ij,t}(\mathbf{x})-\min_{ij,t}(\mathbf{x})}\sqrt{\tfrac{\sum_{ij,t}{(x_{ij,t}-\tilde{x}_{ij,t})^{2}}}{\sum_{ij,t}1}}$ (15)
[…]
$\text{NACRPS}(\hat{F}^{M},\mathbf{x})=\tfrac{\sum_{ij,t}{\text{CRPS}(\hat{F}_{ij,t}^{M},x_{ij,t})}}{\sum_{ij,t}{|x_{ij,t}|}}$ (16)
@@
Baseline Comparison
We now describe our baseline methods. All our experiments were conducted on an HPC cluster using the SLURM scheduler. Each job used up to 256256GB CPU RAM, $8$ CPU cores and one GPU with maximum $40$ GB GPU RAM. Our source code is provided with the supplementary materials.
@@
    DiffSTG (wen2023diffstg): The State-of-the-art (SOTA) diffusion-based model for spatiotemporal forecasting.
@@
    TimeGrad (rasul2021autoregressive) The SOTA diffusion-based model designed for multivariate time-series forecasting.
@@
    DeepVAR (salinas2019high): Combines an RNN-based architecture with a Gaussian copula process output model with a low-rank covariance structure.
@@
    LSTNet (lai2018modeling): Captures local dependency pattern among variables and long-term temporal correlation among time-series trends using CNN and RNN.
@@
    DCRNN (li2017diffusion): A spatiotemporal forecasting framework that combines diffusion convolution and sequence-to-sequence architecture to capture temporal and spatial dependencies.
  6. 6.
    BayesNF (saad2024scalable): SOTA DL method based on a Bayesian Neural Network that maps a multivariate space-time coordinate to a real-valued field.
We used six different patch settings in our experiments as shown in Table  1 (visualization of patches provided in supplementary materials). The experiments are performed at a sparsity level of $95\%$; meaning when forecasting for a patch, only $5\%$ of the cells of that patch will have sensors and inundation values in the context window. During training, we generate random sensor masks ($\mathbf{M}_{k}$) that have $\sim 5\%$ of the cell values to be $1$. For testing, we generate $10$ different sensor masks at random with same ($95\%$) sparsity level for each patch setting. These masks are applied on test data points (from Tidewatch) in a round-robin manner; i.e., mask 1 applied on data point 1, mask 2 applied on data point 2, …, mask 10 applied on datapoint 10, mask 1 applied on datapoint 11, and so on. This ensures that all forecasting methods are being tested with same datapoints. By masking the test datapoints and then making forecasts based on them, we simulate the real-world scenario of forecasting based on sparse sensor observations.
@@
It should to be noted that the number of variables increases quadratically with patch size and linearly with the number of patches. We found that some models do not scale well [^1] We considered $D^{3}$VAE (li2022generative) as a baseline. It scaled poorly; causing memory error for our lowest patch configuration. with increments in the number of variables. Across all experiments, we set the context and prediction length to 12 (experiment results with varying prediction length are provided in the supplementary materials). Eight scenarios were sampled to understand the quality of probabilistic forecasting. Table 1 summarizes the performance of Diff-Sparse with respect to the baseline methods. Some key observations from the results are as follows:
First, we observe that DiffSTG, DCRNN and BayesNF scale poorly to large patch configurations. DiffSTG was unable to process patch configurations where $D>16$. For DCRNN, $(D=64,K=10)$ was the highest patch configuration that was processed successfully, whereas BayesNF fails for $D>64$. In larger settings, we ran out of GPU memory for these methods. It demonstrates the relatively worse computational scalability of these three methods.
Second, in small patch configurations, specifically $D=16\text{ and }K\in\{2,5\}$, Diff-Sparse performs second best. Here, BayesNF is the superior model. Diff-Sparse outperforms the other baselines in these configurations.
Third, in moderate and large patch configurations ($D\in\{64,80,96\}$, $K\in\{10,20\}$), Diff-Sparse outperforms all six baseline methods in terms of both metrics. Across moderate configuration where $D=64$, which is the largest configuration that BayesNF scales to, Diff-Sparse outperforms BayesNF by a margin of $\sim 7$-$24\%$ and $\sim 18$-$29\%$ in terms of NACRPS and NRMSE, respectively. In larger configurations ($D\in\{80,96\},K=20$), Diff-Sparse achieves $59$-$61\%$ improvement over the second-best method (TimeGrad), in terms of the two performance metrics. This illustrates the superiority of Diff-Sparse over baseline methods in performance scalability.